### v2

# Back-Reaction of Super-Hubble Fluctuations, Late Time Tracking and Recent Observational Results

Abstract
Previous studies suggested that the back-reaction of super-Hubble cosmological fluctuations could lead to a dynamical relaxation of the cosmological constant. Moreover, this mechanism appears to be self-regulatory, potentially leading to an oscillatory behavior in the effective dark energy. Such an effect would occur in any cosmological model with super-Hubble matter fluctuations, including the standard $\Lambda$CDM model.
The recent DESI data, which indicate that dark energy may be dynamical, has renewed interest in exploring scenarios leading to such oscillatory behavior. In this work we propose a parametrization to account for the impact of super-Hubble fluctuations on the background energy density of the Universe. We model the total effective cosmological constant as the sum of a constant piece and an oscillating contribution. We perform a preliminary comparison of the background dynamics of this model with the recent radial BAO data from DESI. We also discuss the status of the $H_{0}$ tension problem in this model.
I Introduction
In the standard framework of cosmology, the $\Lambda$CDM model, the dark energy (DE) component, which comprises the majority of today’s cosmological energy density, is parametrized by a positive cosmological constant term, $\Lambda$, in the Einstein equations. Despite its apparent simplicity, this interpretation faces conceptual problems. In addition to some well known problems, such as the cosmological constant problem, the coincidence problem and the Hubble tension problem, several works in the literature have been pointing to an instability of a de Sitter phase. In particular, there are indications that a de Sitter space could be unstable due to infrared (IR) effects Polyakov:2012uc ; Polyakov:2007mm ; Mazur:1986et ; Mottola:1985qt ; Mottola:1984ar ; Tsamis:1994ca  [^1] In the context of String Theory there have also been indications that it is not possible to obtain a stable de Sitter phase Obied:2018sgi ; Dvali:2014gua ; Danielsson:2018ztv ; Dasgupta:2018rtp , and the recently suggested “Trans-Planckian Censorship Criterion” Bedroya:2019snp ; Bedroya:2019tba ; Brandenberger:2021pzy implies that a positive cosmological constant is not consistent from an effective field theory point of view.. Already in the work of Ref. Tsamis:1992sx ; Tsamis:1996qq it has been suggested that the reason that our universe is not inflating today is that IR processes in pure quantum gravity (no matter) tend to screen the bare $\Lambda$. However, the authors of Tsamis:1996qq found that these effects become important only after perturbation theory has broken down. In 1997, the work of Ref.Abramo:1997hu studied the back-reaction of super-Hubble modes of cosmological perturbations induced by matter, finding that the effective energy density associated with these modes would counteract any pre-existing $\Lambda$ (see also Finelli:2001bn ; Finelli:2003bp ; Marozzi:2006ky ). Later, in 1999, in the work of Brandenberger:1999su it was speculated that this dynamical relaxation mechanism for $\Lambda$ could be self-regulating  [^2] Note also the related “Everpresent Lambda” scenario, which was proposed by drawing on ideas from causal set theory and unimodular gravity, and which leads to an effective fluctuating cosmological “constant”, see Refs.Ahmed:2002mj ; Das:2023hbw ; Das:2023rvg , for instance.. A crucial question that remained at that time was whether this back-reaction effect could be locally measured Unruh:1998ic ; Geshnizjani:2002wp ; Abramo:2001dc . Recent works have shown that super-Hubble fluctuation modes can in fact modify the parameters of a local Friedmann cosmology, when measured by a clockfield Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr ; Comeau:2023gxk ; Comeau:2023euf  [^3] In the context of the back-reaction of sub-Hubble cosmological fluctuations, a recent analysis can be found, for instance, in Ref.Giani:2024nnv  [^4] Another interesting approaches to account for the back-reaction effect can be seen in the works of Refs.Kolb:2009rp ; Kolb:2011zz ; Marra:2011ct , for instance.. In the work of Brandenberger:2018fdd the analysis was extended beyond perturbation theory, showing that the locally measured Hubble expansion rate indeed obtains a negative contribution from the back-reaction of super-Hubble fluctuations, whose amplitude grows in time. This supports the suggestion that (quasi) de Sitter space-times are unstable, and that this back-reaction mechanism might lead to a dynamical relaxation of the cosmological constant. As first speculated in Ref.Brandenberger:1999su , this is expected to lead to an effective dark energy density which oscillates in time around the dominant matter density.
The quest for understanding the nature and dynamics of the accelerated expansion of the Universe became even more interesting recently with the improvement in the measurements of the expansion history, which are increasingly allowing us to constrain the expansion history of the Universe. Recently, the Dark Energy Spectroscopic Instrument (DESI) Year 1 (DR1)DESI:2024aqx ; DESI:2024kob ; DESI:2024mwx and DESI DR2 DESI:2025zgx ; DESI:2025fii results were released, officially marking the start of the Stage IV DE era. The cosmological results stem from measurements of baryon acoustic oscillations (BAO) in galaxy, quasar and Lyman-$\alpha$ forest tracers. The DESI results, when combined with a number of external probes, provide intriguing hints for a dynamical, time-evolving DE component DESI:2024mwx ; DESI:2024aqx ; DESI:2024kob . These hints were present already in the DR1 release  [^5] See, however, Payeur:2024dnq for caveats.. Later, the enhanced statistical power and extended redshift coverage of DESI DR2 sustained the preference for a dynamical dark energy component, yielding a level of statistical significance comparable to, or greater than, that reported with DESI DR1 DESI:2025wyn ; DESI:2025fii ; DESI:2025zgx . These hints may have enormous implications for the understanding of the nature of DE, and align with expectations from certain fundamental theory frameworks Brandenberger:2025hof .
In light of these results, recent works, among them Jiang:2024xnu and Escamilla:2024fzq , have pointed to signals, from late-time cosmological data, of oscillatory/non-monotonic features in the shape of the expansion rate at low redshifts  [^6] See, however, Akarsu:2022lhx for a warning that oscillatory features in $H(z)$ can be induced when fitting for this function given incomplete data. Also, the DESI full shape modelling does not show a signal Colgain:2024mtg .. The possibility of an oscillating dark energy has already been addressed in earlier works, as for instance in Refs. Zhao:2017cud ; Tamayo:2019gqj ; Rubano:2003er ; Linder:2005dw ; Feng:2004ff ; Nojiri:2006ww ; Kurek:2007bu ; Jain:2007fa ; Saez-Gomez:2008mkj ; Kurek:2008qt ; Pace:2011kb ; Pan:2017zoh ; Panotopoulos:2018sso ; Rezaei:2019roe ; Yao:2022jrw ; Rezaei:2024vtg ; Tian:2019enx ; DESI:2024kob ; Adil:2023exv ; Mbonye:2022cnf ; Mbonye:2024mss . In Ref. Escamilla:2024fzq , in light of the recent cosmological data, including DR1, twelve different parametrizations for an oscillatory dark energy equation of state were analyzed, among them the ones proposed in Refs. Linder:2005dw ; Feng:2004ff ; Pan:2017zoh ; Zhao:2005vj . This analysis demonstrated that all the oscillating DE parametrizations considered provide a better fit to the DR1 when compared to the cosmological constant, although in the Bayesian evidence analysis they are penalized due to the additional free parameter. In a similar context, in the work of Ref. Jiang:2024xnu , the late-time expansion history was reconstructed in a way, according to the authors, to be “as model-independent and non-parametric as possible” (see also Heisenberg:2020ywd ; Haude:2019qms ). In their first reconstruction, using DR1 + PantheonPlus data, two features in the unnormalized expansion rate were identified, a bump at redshift $z\sim 0.5$, and a depression at $z\sim 0.9$ (both features were absent when replacing the DESI dataset with the older SDSS one). On the other hand, the depression feature was noticeable in both the DR1+PantheonPlus and DR1+Union3 reconstructions, but not in the DR1+DESY5 one  [^7] For more information on the reconstruction methods see Ref.Jiang:2024xnu .. While the Chevallier-Polarski-Linder (CPL) fit, the $w_{0}-w_{a}$ parametrization often considered when interpreting DESI data, is able to partially capture the data characteristic behaviour, it cannot fit all the features, since the model does not have sufficient complexity. Achieving this may require an oscillatory/non-monotonic behavior of the Hubble expansion rate Jiang:2024xnu . More recently, some works (see Kessler:2025kju for instance) found evidence for an oscillating behaviour of the dark energy equation of state also from the more recent DR2 data.
While a more conclusive result will need to wait for more precise future data, we believe it is now important to further investigate theoretical mechanisms that could account for a non-monotonic dynamical behavior of DE. As discussed above, the back-reaction effect of super-Hubble fluctuations is one mechanism that could naturally lead to an oscillatory behavior for the dark energy density, and consequently for the Hubble parameter, without the need for extra ingredients beyond the standard $\Lambda$CDM model. This back-reaction effect is present in any cosmological model whenever super-Hubble cosmological fluctuations are present.
In the present work we propose a parametrization to account for the back-reaction effect of super-Hubble fluctuations by means of an oscillating effective energy density component. This may provide a full non-perturbative description of the back-reaction effect. We then analyze such a model considering data from the radial component of the Baryon Acoustic Oscillations from the recent DESI DR2 data. We compute the Hubble parameter at low redshifts as predicted by this model, and compare the result with the data in order to place constraints on the maximum allowed amplitude of the oscillations. We also analyze the predictions of the model for the effective dark energy equation of state, for the deceleration parameter, and for $H_{0}$.
Our study can be considered to be a first-step analysis limited to background quantities, aiming only to highlight the importance of not neglecting the back-reaction mechanism. Nevertheless, in order to obtain more conclusive results, it is necessary to further analyze the predictions of the model, in particular taking into account the induced cosmological perturbations. This will allow us to test the model using additional BAO data beyond only the radial component, as well as using further data from other cosmological sources. This analysis is left for a future work.
The paper is organized as follows: In Section 2 we propose a new parametrization to describe the back-reaction of super-Hubble fluctuations in the late universe. In Section 3 we discuss the data analysis methods which we use to compare our model predictions with the recent DESI results. In Section 4 we present our results and we conclude in Section 5.
II A Model for the Back-Reaction of Super-Hubble fluctuations
The Einstein field equations, as its well known, are highly nonlinear. Even at a classical level fluctuations at second order affect the background. This effect is called back-reaction. Matter fluctuations induce metric fluctuations through the Einstein equations. These metric fluctuations also back-react on the homogeneous background metric. The full evolution of the metric is governed by the Einstein equations,
[…]
where $G_{\mu\nu}$ is the Einstein tensor of the metric $g_{\mu\nu}$, $G$ is Newton’s gravitational constant and $T_{\mu\nu}$ is the energy-momentum tensor of matter.
Metric and matter (taken for simplicity to be a scalar matter field $\varphi$) can be written, at linear order in the amplitude of cosmological fluctuations, as
$ds^{2}\,=\,(1+2\phi(x,t))dt^{2}-a(t)^{2}[1-2\psi(x,t)\gamma_{ij}dx^{i}dx^{j}],$ (2)
and
$\varphi(x,t)\,=\,\varphi_{0}+\delta\varphi(x,t).$ (3)
In the absence of anisotropic stress, the scalar metric fluctuation variables $\phi$ and $\psi$ in eq.2 are equal. To write the metric, we have chosen a particular coordinate system (longitudinal gauge)  [^8] See e.g. Mukhanov:1990me for a review of the theory of cosmological perturbations.. In Eq. 2 $\gamma_{ij}$ is the background metric of the constant time hypersurfaces. We assume vanishing spatial curvature and then $\gamma_{ij}=\delta_{ij}$. In the above equations, the background metric is given by the scale factor $a(t)$ and the background matter by $\varphi_{0}(t)$. We will only consider here the effects of scalar metric fluctuations.
In order to see how the linear fluctuations affect the background, we substitute the ansatz 2 and 3 into the Einstein equations and expand to second order. The zero’th order terms satisfy the background equations. The linear terms are assumed to obey the linear perturbation equations. In this case, the Einstein equations are not satisfied at quadratic order. Back-reaction is a way to correct for this at quadratic order. The quadratic terms on the left hand side of 1 can be moved to the right hand side of the equation, where they form an effective energy-momentum tensor when combined with the quadratic terms in $T_{\mu\nu}$ Abramo:1997hu ; Mukhanov:1996ak ,
$\tau_{\mu\nu}\;\equiv\;<T_{\mu\nu}^{(2)}-\frac{1}{8\pi G}G_{\mu\nu}^{(2)}>,$ (4)
where the superscript (2) indicates the order of the terms.
At second order, the linear perturbations modify the background metric. This effect can be described by introducing a modified background metric
$g_{\mu\nu}^{bg}(t)\,\equiv\,g_{\mu\nu}^{0}(t)+\delta g^{bg,2}(t)\,,$ (5)
where the first term is the original background metric and the second reflects the quadratic corrections to the background.
By taking the spatial average of the Einstein equation 1, expanded to second order, we can describe the corrections to the background. Following the method proposed in Abramo:1997hu ; Mukhanov:1996ak (and reviewed in Brandenberger:2002sk ) we can write the resulting equation in the form,
[…]
where the first term on the right hand side of the equation is the contribution of the background matter field to $T_{\mu\nu}$. Improved studies of the back-reaction of super-Hubble fluctuations focus on physically measurable quantities such as the spatial average of the local Hubble expansion rate Geshnizjani:2003cn ; Marozzi:2012tp ; Finelli:2011cw ; Gasperini:2009wp ; Gasperini:2009mu ; Marozzi:2010qz ; Marozzi:2011zb ; Abramo:2001dd ; Losic:2005vg ; Losic:2006ht ; Brandenberger:2004ix ; Afshordi:2000nr even beyond perturbation theory Brandenberger:2018fdd . Some of these results have pointed to a self adjusting mechanism for the cosmological constant, as first discussed in Brandenberger:1999su .
The origin of the self-adjustment mechanism for the relaxation of the cosmological constant is as follows Brandenberger:1999su .: in an accelerating phase, fluctuation modes exit the Hubble radius and freeze out. Cosmological perturbations induced by super-Hubble matter fluctuations contribute to the effective energy-momentum tensor which has the form of a negative cosmological constant (negligible kinetic and tension energies leading to an effective equation of state $w=-1$, and negative effective energy density - negative since matter induces a potential well and the negativity of the gravitational energy overwhelms the positivity of the matter energy for super-Hubble modes). This will induce a decrease in the effective cosmological constant by counteracting the effect of the pre-existing $\Lambda$. But once the energy density $\rho_{DE}(t)$ in the effective cosmological constant has fallen below that of matter, accelerated expansion will stop, modes will re-enter the Hubble radius, the decline in $\rho_{DE}(t)$ halts, and within a fraction of the Hubble time, a renewed buildup of the backreaction effect takes place.
In light of this, we propose here the following parametrization for the effective energy density associated to the back-reaction of super-Hubble fluctuations:
$\rho_{BR}\ =\,A(t)\;cos(\omega t+\phi)\;\rho_{m}(t),$ (7)
where $A$ is a dimensionless amplitude, $\omega$ is the frequency of the oscillation, $\phi$ is a phase and $\rho_{m}$ is the background density of matter. The term $\omega t$ in the above equation can be written in terms of the redshift $z$ as $\omega t\equiv nz$, with $n$ being a constant, which we expect to have an order of magnitude not much smaller or bigger than one, since the typical time scale of oscillations is expected to be set by the Hubble time. This ansatz expresses the idea that, after the time $t_{eq}$ of equal matter and radiation, the back-reaction term should track the energy density $\rho_{m}$ of matter 999 Note that an oscillatory contribution to the dark energy density was recently also shown to emerge from a model in which the dark energy field emerges from the dynamics of a highly curved field space motivated by string theory Payeur:2024kyy .
A priori it is expected that the amplitude $A(t)$ should be time-dependent. However, in the present work, we will focus on the late time predictions, since we are going to analyze our model in light of data in the range $2.50\leq z<$ 2.5. For this small redshift interval we believe it is reasonable to consider $A(t)$ to be constant. Therefore, in the rest of the paper we are going to consider the following expression for the effective density from the back-reaction effect
$\rho_{BR}\ =\,A\;cos(nz+\phi)\;\rho_{m}(t)\,.$ (8)
The contribution of the energy density from the back-reaction effect is added to the total energy of the universe, in such a way that the first Friedmann equation can be written as,
[…]
where we have kept a bare cosmological constant term $\Lambda$. We will set the speed of light $c$ to one in the following.
By substituting Eq.7 in the above equation, we can write the first Friedmann equation as follows,
$H^{2}\,=\,H_{0}^{2}\;\left((1+A\;cos(nz+\phi))\;\Omega_{m0}\;(1+z)^{3}+\Omega_{\Lambda 0}\right),$ (10)
where  [^10] The subscripts 00 indicate the quantities at the present time.
[…]
Another important quantity to be analyzed is the deceleration parameter, which must show a transition from a decelerated regime to the accelerated one as the dark energy starts dominating. The deceleration parameter is $q\equiv H^{\prime}/(aH)-1$ where the prime indicates derivative with respect to the redshift. By using the above equation, we obtain for the deceleration parameter the following expression,
[…]
Concerning the phase $\phi$, we will choose it in a way that the back-reaction mechanism begins at a redshift $z=3400$ with $\rho_{BR}=0$, when it starts building up. This choice is reasonable since we are taking the back-reaction energy density as being proportional to the matter energy density, and before recombination it would rather track the radiation component.
We can now compute the equation of state associated to the back-reaction effective fluid. The continuity equation is not valid in the usual form for this fluid, since there is an energy transfer from the cosmological fluctuations to this effective component Abramo:1997uy . In this case, in order to obtain $w_{BR}$, we can consider the second Friedmann equation for the total fluid (considering all the components of the Universe),
$\dot{H}+H^{2}=-\frac{4\pi G}{3}\left(\rho_{tot}+\frac{3p_{tot}}{c^{2}}\right)=-\frac{4\pi G}{3}\left(\frac{3H^{2}}{8\pi G}-\frac{3\rho_{\Lambda}}{c^{2}}+\frac{3w_{BR}\;\rho_{BR}}{c^{2}}\right).$ (13)
By isolating $w_{BR}$ in the above equation and considering that $\dot{H}=-H^{\prime}H/a$, we obtain the following expression for the equation of state,
$\begin{split}w_{BR}&=\left(\frac{H^{\prime}H}{a}-\frac{3H^{2}}{2}+\frac{3H_{0}^{2}\;\Omega_{\Lambda 0}}{2}\right)\frac{1}{4\pi G}\;\frac{1}{\rho_{BR}},\\ &=-\frac{n\left(z+1\right)}{3\tan\left(n(z-3400)\right)}\end{split}$ (14)
In Section IV we will illustrate the behavior of the Hubble parameter at low redshifts, the deceleration parameter and the effective equation of state for chosen values of the frequency and amplitude in Eq.7.
In this work we will restrict our analysis to the background dynamics. We then do not need to make extra assumptions on the cosmological model at perturbative level. A further discussion of the perturbative quantities of the model is left for a future work. While this allow us to make more general predictions concerning the back-reaction mechanism, we consider this study to be a preliminary analysis. In order to obtain more conclusive results, it would be necessary to go beyond the background analysis and test the model with a more complete cosmological dataset. However the expression above already provide us some insights in light of specific datasets, as for instance the radial component of BAO, which we are going to discuss in the next section.
III Methods and Data Analysis
The baryon acoustic oscillation (BAO) method is one of the main methods already developed to measure the expansion history of the Universe (for a more complete discussion see for instance Weinberg:2013agg ; Chen:2024tfp ; DESI:2024uvr ). In the pre-recombination Universe, acoustic oscillations in the baryon-photon fluid imprinted a characteristic scale on the clustering of matter, which manifests itself as a bump in the two-point correlation function of matter (or as a series of oscillations in the power spectrum). The comoving scale of this feature, is given by the sound horizon at the end of the baryon drag epoch, $r_{d}\approx 100h^{-1}Mpc$. This quantity depends on the photon and baryon content of the universe at this epoch, and it is well constrained by CMB measurements. The BAO feature can be traced by galaxies, quasars, and the Lyman-$\alpha$ forest.
The BAO signature can be measured in two ways: one can use the angular diameter distance to determine the physical length scale corresponding to a certain angle on the sky of the observer, or one can use the difference in redshift to infer some physical radial distance. Therefore, there are measurements of the apparent size of the BAO standard ruler perpendicular and parallel to the line of sight, allowing one to constrain the angular diameter distance $D_{A}(z)$ and the Hubble parameter $H(z)$, respectively. If the Hubble parameter of our Universe is actually different than the one predicted by the model considered, then there will be differences in the radial BAO scale.
We can therefore decompose the BAO locations into transverse and line-of-sight dilation parameters, which are denoted by the Alcock-Paczynski-like dilation parameters $\alpha_{\perp}$ and $\alpha_{||}$, which are linked to the angular diameter distance, and to the Hubble parameter, at a given redshift. Here in this work we are interested in the line-of-sight parameter, which is defined as DESI:2024uvr ,
$\alpha_{||}\,\equiv\,\frac{H^{fid}(z)\;r_{d}^{fid}}{H(z)\;r_{d}},$ (15)
where the “fid” denotes the quantities measured in the fiducial cosmology, which is the Planck-normalized $\Lambda$CDM model used to convert redshifts into distances in the survey  [^11] Using the notation of Ref.DESI:2025zgx , we can write $\alpha_{||}=\alpha_{\text{iso}}\alpha_{\text{AP}}^{2/3}$.. In the above, $r_{d}$ is the comoving sound horizon evaluated at the redshift at which baryon-drag optical depth is equal to unity, and it depends on quantities at redshifts higher than that. In the case of the back-reaction model presented in Section II, the quantity $r_{d}$ will be considered to be equal to the fiducial one, since we are here not investigating the effects of back-reaction at earlier times. Therefore in our analysis we are going to consider the approximate expression [^12] For a similar analysis in a different context, see for instance the work of RefClifton:2024mdy ,
$\alpha_{||}\,\approx\,\frac{H^{fid}(z)}{H(z)}\implies\frac{H(z)-H^{fid}(z)}{H^{fid}(z)}=\frac{\Delta H(z)}{H^{fid}(z)}=\frac{1}{\alpha_{||}}-1,$ (16)
where $\Delta H\equiv H(z)-H^{fid}(z)$. We are going to use information from the DESI DR2 galaxy, quasar and Lyman-$\alpha$ results DESI:2025zgx to interpret the constraints on $\alpha_{||}$ in terms of constraints on $\Delta H$ at given redshifts.
The data we are going to consider are described in the following table:
Tracer Redshift $\alpha_{||}$ $\Delta H/H^{fid}$
LRG1 $0.51$ $0.9615419512\pm 0.01922057331$ $0.03999622561\pm 0.0207888212$
LRG2 $0.706$ $0.9647661889\pm 0.01642269702$ $0.03652057001\pm 0.01764413334$
ELG2 $1.321$ $0,9932025202\pm 0.01256687344$ $0.006844001718\pm 0,01273947749$
QSO $1.484$ $1.02675147\pm 0.02099900114$ $-0.02605447469\pm 0.01991902012$
Lya $2.33$ $1.008212343\pm 0.01700856044$ $-0.008145449517\pm 0.01673260419$
LRG3 $0.922$ $0.9953917692\pm 0.04052811979$ $0.004629564906\pm 0.04090424355$
ELG1 $0.955$ $1.001814039\pm 0.01652471677$ $-0,001810754569\pm 0,01646492654$
Table 1: Dataset used for $\alpha_{||}$ from the DESI DR2 galaxy, quasar and Lyman-$\alpha$ results taken from Ref.DESI:2025zgx , and its interpretation in terms of $\Delta H/H^{fid}$ at given redshifts.
Our procedure is the following: We first solve Eq.10 (with $\Lambda_{0}$ taken to be the same as in the fiducial cosmology) in order to obtain H(z) predicted by the back-reaction model. From this we can plot the resulting function $\Delta H/H_{\Lambda CDM}$ for the model together with the observational values and error bars shown in Tab.1. However, in order to plot the quantity $\Delta H/H_{\Lambda CDM}$ predicted by the back-reaction model, the parameters in equation 7 must be fixed. As mentioned earlier, the phase is fixed in such a way that $\rho_{BR}$ is zero at redshift $z=3.400$. The amplitude $A$ and the ”frequency”, $nz$, remain to be fixed. We will focus on values of the amplitude which are of the order of $A\sim 10^{-1}$. Larger values of the amplitude would give rise to too large oscillations of the Hubble parameter, while smaller values would not lead to observable differences compared to our fiducial cosmology. We cannot exactly determine the amplitude from the considerations in Section II since this would require a non-perturbative calculation that has not yet been achieved. On the other hand, the considerations in Section II imply that the frequency of oscillation is set by a fraction of the Hubble rate. Hence, we would expect that a non fine-tuned frequency would correspond to a value of $n$ with order of magnitude approximately equal to one.
We show the results obtained in the next Section.
IV Results
In this section we show the predictions of the back-reaction model compared with the data of Ref.DESI:2025zgx . In Fig. 1 we show the late-time behavior of the quantity $\Delta H/H_{\Lambda\text{CDM}}$ (curves in color) as predicted by the back-reaction model, for values of the amplitude ranging from zero (which represents $\Lambda\text{CDM}$) to $A=-0.15$ and for a fixed frequency characterized by the parameter $n\approx 2.4$ introduced in Eq.10. These predictions are compared with the DESI DR2 results (red dots and error bars). Different shapes of the theoretical curve can be obtained using different values of the frequency and amplitude. We choose, just for the sake of illustration, examples of values that provide a shape for the curves which tracks the DESI data better than the standard $\Lambda$CDM model. In addition, in choosing the frequency value to be illustrated in the Figures, we selected models which can alleviate the Hubble tension even slightly, and we disregard models with too high frequencies ($n\gtrsim 5$) since they could lead to smaller $\chi^{2}$ simply due to overfitting. Among the frequency values that satisfy these criteria, we chose the one that provides the smaller $\chi^{2}$, which corresponds to $n\approx 2.4$. However, higher values for the frequency can be considered, also providing a $\chi^{2}$ value lower than the $\Lambda$CDM. Fig.6 in the Appendix shows the same plot as Fig.1 but for DR1 data. Comparing Fig.1 with Fig.6 we see that the evidence of an oscillating dark energy component is stronger in DR2 than in DR1.
Note that the values of the free parameters (amplitude and frequency) which we have used are within the range of values that can be considered as “reasonable” from the theoretical point of view. In particular, the frequency is in the range expected from qualitative arguments.
Figure 1: $\Delta H/H$ predicted by the back-reaction model compared with DESI DR2 data
Although the results look promising, our preliminary analysis should not be viewed as conclusive in terms of statistical evidence. With future data releases, the error bars are expected to shrink, and by combining those with datasets from other sources it should then be possible to obtain more solid conclusion on the preference of the data for $\Lambda$CDM or a dynamical dark energy. The objective of Fig. 1 is to show the range of amplitudes of the oscillating dark energy term which can be in agreement with the DESI constraints. We can see that, for the chosen value $n\sim 2.4$, values as low as $|A|\sim 0.05$ can fit the data within the observational uncertainties.
As a preliminary analysis, we performed a test minimizing the $\chi^{2}$ of the model with respect to the DESI DR2 data. The method used consists of executing a set of minimization routines, such that for each minimization sequence we chose the initial values of $A$ to lie within the range $[-0.1,0.1]$, and those of $n$ within $[3,5]$ These intervals correspond to initial guesses of suitable parameter values and are starting points for the parameter exploration.
The expression to be minimized is the $\chi^{2}$ of $\Delta H(z)/H_{\Lambda CDM}$ (which expresses the relative difference between the proposed model and the standard cosmological model as a function of redshift) with respect to the experimental data from DESI DR2. As described previously, the DESI experimental results were interpreted in terms of $\Delta H$, and the error propagation was calculated. Some minimization results lead to a shape more in line with the one indicated by DESI data than others. However, in practically all the cases considered, for the mentioned initial values, the $\chi^{2}$ obtained were smaller than the one of the $\Lambda$CDM model. On the other hand, these analysis were restricted to amplitudes with order of magnitude $A=\mathcal{O}(10^{-1})$ or smaller. For larger values of $A$ the fit is poor. This is shown in Fig.2, where the value of $\chi^{2}$ is plotted as a function of the amplitude. From this same figure we can also see that for values of the amplitude higher than zero and lower than $A\sim-0.15$ the back-reaction model provides a fit worse than the standard model (which is recovered for the value $A=0$, the star in the plot). We can also see from Fig. 2 that the minimum value of the $\chi^{2}$ is shifted with respect to the $\Lambda$CDM value. A similar plot using the data from DR1 can be seen in the Appendix. By comparing the $\chi^{2}$ as a function of the amplitude for the two datasets, DR1 and DR2, we see that the minimum value of the $\chi^{2}$ does not change considerably. However from the overall analysis, we obtain that the frequencies that tend to provided best fitting for the DR2 data are in general smaller then the ones that fits better the DR1 data. This general tendency seems to be in agreement with the results of Ref. Ormondroyd:2025iaf . While the error bars still do not allow us to conclude that our model performs better than $\Lambda$CDM, these results can be viewed as an indication of the importance of further studies of this model, rather than a firm statistical conclusion.
Figure 2: The value of $\chi^{2}$ for the back-reaction model with respect to DESI data as a function of the parameter $A$. The value for $\Lambda$CDM is represented by a star.
In Figs 3 and 4 we plot, respectively, the deceleration parameter and the equations of state (EOS) of the back-reaction model (the EOS for the effective backreaction fluid and the total EOS including all the components), according to the equations in Section II. The pair of values $(A\sim-0.08,n\sim 2.4)$ was chosen as an illustrative example in all these figures. We can see in Fig.3 that this model is able to correctly predict the transition from decelerated to accelerated expansion around the expected redshift.
Figure 3: Deceleration parameter from $z=2.5$ to $z=0$ predicted by the back-reaction model, together with the $\Lambda$CDM prediction.
In Fig.4 (left) we illustrate the evolution of the equation of state for the effective back-reaction fluid, given by Eq.14. We considered again the same values for the parameters $A$ and $n$. We can see in this plot that the equation of state shows the average value $w\sim 0$. One can also see that, at the exact redshifts when $\rho_{BR}$ crosses zero, a spike can be seen in the plot of $w_{BR}$. This should be expected from the fluid equation of state $w_{BR}\equiv p_{BR}/\rho_{BR}$. The spikes correspond to the points where the quantity $\Delta H/H_{\Lambda CDM}$ crosses zero in Fig.1. However, this behavior does not correspond to a problem  [^13] A similar behavior can be seen for instance, in the model analyzed in Ref.Tiwari:2024gzo . since the observable quantities in cosmology correspond to integrals of the equation of state over a period of time. Therefore these spikes are not expected to be present in any cosmological observable. In addition, the contribution from the back-reaction effective fluid is always subdominant compared to the other components of the Universe. The effective equation of state of the total fluid has the expected overall behavior, as shown in Fig.4 (right), i.e., it oscillates around the $\Lambda CDM$ prediction.
Figure 4: The evolution of the effective equation of state of the back-reaction fluid (left) and the total effective equation of state (right).
Another major challenge that has become more pressing following the DESI data is the Hubble tension problem Colgain:2025nzf ; Poulin:2025nfb ; Pang:2025lvh . Recent results from DR2 bolster the case for a dynamical dark energy scenario exhibiting phantom crossing behavior. This framework yields a lower inferred value of the Hubble constant, thereby intensifying the Hubble tension — which can be regarded as a symptom of potential underlying inconsistencies in the standard model. Therefore it is interesting to analyse in which cases, for which values of the parameters, the back-reaction model considered can provide some relief to the Hubble tension. In order to shed some light to this question we perform an analysis of the value of $H_{0}$ predicted for different combinations of the parameters $n$ and $A$. The analysis uses the method described previously. We show in Fig. 5 a plot of $H_{0}$ ($H_{0}=H_{0}(A,n_{\text{best}})$) as a function of $A$, where $n_{\text{best}}\sim 2.4$ is the value of $n$ chosen from the minimization method performed. We can see in the plot that, for the representative chosen values of the parameters, the $H_{0}$ predicted by our model is very close to the $\Lambda$CDM one. This is partly due to the choice of the fixed A and n and also due to the fact that the back-reaction contribution is proportional to $\rho_{m}$, therefore decreasing at later times, causing the difference of our model with respect to $\Lambda$CDM also to decrease.
Figure 5: $H_{0}$ as a function of $A$, for $n$ chosen from the minimization method performed.
From all the results of our minimizations, the highest value that was obtained was $H_{0}\approx 1.009\;H_{0\Lambda CDM}$, which is achieved for $A\approx-0.07$ and $n\approx 2.06$. For the amplitude $A=-0.05$, for instance, we obtain $H_{0}/H_{0\Lambda CDM}=1\pm 10^{-3}$, i.e., it’s basically indistinguishable from $\Lambda CDM$. Nevertheless, it could be possible to solve the Hubble tension given a sufficiently high absolute values of the amplitude. However, from Fig 1 we can see that such high absolute values of $A$ do not provide a good fit to DESI data. If we extrapolate and investigate the Hubble constant value for an amplitude as low as $A=-1$, we obtain $H_{0}\approx 1.09\;H_{0\Lambda CDM}$, considering $n=n_{\text{best}}$.
It is important to emphasize again that a priori there is no reason that the amplitude $A$ should be independent of time. In fact, from the considerations of back-reaction it is not unreasonable to expect that the effective amplitude of the oscillating contribution to DE performs damped oscillations. In this case, $A$ could be large at early times and then slowly decay to a value consistent with the current late time data. This extended scenario might therefore be able to provide a solution of the Hubble tension. On the other hand, without a careful analysis of the fluctuations produced in the model, it is not possible to gauge whether such scenario would be consistent with other constraints, e.g. the value of $\sigma_{8}$.
While conclusive results concerning the viability of the model will need to wait for a more complete analysis with the future data, our results shows that the inclusion of back-reaction effects from super-Hubble fluctuations can offer valuable insights in light of recent observational data.
V Conclusion
In the present work, we studied a toy model of oscillating dark energy motivated by considerations of the back-reaction of super-Hubble fluctuations in the late universe. We proposed a parametrization of the oscillating dark energy component to account for the back-reaction effect on the background cosmology (based on the conjectured dynamical scaling fixed point solution describing the back-reaction effect at a fully non-perturbative level). We computed the Hubble parameter at low redshifts predicted by the model, obtaining an oscillatory dynamics. We also compared the resulting background cosmology with data from the radial component of the Baryon Acoustic Oscillations obtained from the recent DESI survey DESI:2025zgx . We found that, for a frequency scale expected from theoretical considerations, a fit to the recent observations can be obtained for a certain range of the amplitude of the oscillation, which is a free parameter in the analysis, and for a frequency of oscillation motivated by the theory. At the present time the observational error bars are still too large to observe a statistically significant advantage of our model over the standard $\Lambda$CDM cosmology. But, with upcoming observations the error bars will shrink and may lead to a way to extract a non-vanishing amplitude of an oscillating component to dark energy in a reliable way.
We also investigated the impact of our model on the Hubble tension. We found that for a large amplitude $A$ of the oscillatory component (such that the energy density in the oscillating term is comparable to the matter energy density) a sufficiently large change in $H_{0}$ results to explain the Hubble tension. However such a high amplitude does not provide a good fit to the data. If, on the other hand, $A$ depended on time and decreased from a large value near the time of recombination to a small value at the present time, the Hubble constant tension might be successfully addressable. However, detailed studies of the effects of our model on cosmological fluctuations is required to check the consistency of such a scenario.
The next step should in fact be to analyze the perturbations in our scenario. This would allow us to make comparisons with other observable quantities, e.g. CMB anisotropies, and in particular the Integrated Sachs Wolfe (ISW) effect. In this context, in Ref.Pace:2011kb the behavior of several cosmological observables, such as the linear growth factor, the ISW effect, the number counts of massive structures, and the matter and cosmic shear power spectra were estimated for several oscillating dark energy models. It was shown that, for several of the models considered, independently of the amplitude and the frequency of the dark energy oscillations, none of these observables showed a significant oscillating behavior as a function of redshift. This is a consequence of the fact that such observables are integrals over functions of the expansion rate along the cosmic history, consequently smoothing oscillatory features below the level of detectability. This analysis was performed under certain assumptions, for instance, supposing that dark energy does not cluster on the scales of interest. In addition, the models studied in that reference have a different behavior compared to the back-reaction model we consider in our work. A complete analysis specific for our scenario is hence still missing, but we believe it might be possible that the main conclusions of the work of Pace:2011kb might also hold for our model. As the observables mentioned are integrals over functions of the expansion rate along the cosmic history, we expect that also in our scenario the oscillatory features will be smoothed, keeping the ISW effect in agreement with current observations.
Another quantity that is going to be investigated in upcoming work is the sound speed predicted in our scenario. Although it may not necessarily be true that the back-reaction model must reproduce the $\Lambda$CDM behavior for this quantity, one could expect similar predictions in this regard. Since our model accounts only for the back-reaction of super-Hubble fluctuations, one could expect that a quantity related to the dynamics of sub-Hubble fluctuations, as the sound speed, could follow a similar behavior as the standard model. A further analysis of these aspects will be addressed in a future work.
Our work can be considered to be a first-step analysis limited to the background predictions of the model. Our results, although preliminary, highlights the importance of considering the back-reaction effect of super-Hubble fluctuations in the late Universe, as such back-reaction effects are expected to be present in any realistic scenario. We have shown that this provides a dynamics which is interesting in light of recent observational results.
Appendix A Previous DESI DR1 Analysis
In this Appendix we show the analysis of the back-reaction model in light of the previous DR1 results for the sake of comparison. Using the DR1 data we produce the Table below with the values for $\alpha_{||}$ from Refs. DESI:2024uvr DESI:2024lzq , and its interpretation in terms of $\Delta H/H^{fid}$ at given redshifts.
Tracer Redshift $\alpha_{||}$ $\Delta H/H^{fid}$
LRG 1 $0.4-0.6$ $0.9219\pm 0.0269$ $0.0847\pm 0.0317$
LRG 2 $0.6-0.8$ $0.9955\pm 0.0296$ $0.0045\pm 0.0299$
LRG 3 $0.8-1.1$ $1.0057\pm 0.0214$ $-0.0057\pm 0.0212$
ELG 2 $1.1-1.6$ $0.9797\pm 0.0297$ $0.0207\pm 0.0309$
Lyman-$\alpha$ $\sim 2.33$ $0.993^{+0.029}_{-0.032}$ $0.007^{+0.029}_{-0.032}$
Table 2: Dataset used for $\alpha_{||}$ from the DESI first-year galaxy and quasar results DESI:2024uvr , and the Lyman-$\alpha$ forest DESI:2024lzq , and its interpretation in terms of $\Delta H/H^{fid}$ at given redshifts.
In Fig.6 we show the late-time behavior of the function $\Delta H/H_{\Lambda CDM}$ (curves in color) as predicted by the back-reaction model, for values of the parameter A ranging from zero to $A=-0.15$ and for a fixed value of $n\sim 5$. These predictions are compared with the DESI DR1 results (red dots and error bars). Different shapes of the theoretical curve can be obtained using different values of the frequency and amplitude. To illustrate the method, we select particular sets of parameter values that result in curve shapes that more closely align with the DESI DR1 data.
Figure 6: $\Delta H/H$ predicted by the back-reaction model compared with DR1 data. In all the curves we considered the value $n\sim 5$.
We performed a test minimizing the $\chi^{2}$ of the model with respect to the DR1 data using the same method described in Section IV. In Fig.7, we show the value of $\chi^{2}$ plotted as a function of the amplitude. As in the analysis for DR2 shown in Fig.1, we can see that also in this case higher values of the amplitude the back-reaction model provides a fit worse than the standard model. Again we can see that the minimum value of the $\chi^{2}$ is shifted with respect to the $\Lambda$CDM value (which is recovered for the value A = 0). The minimum is achieved for an amplitude slightly smaller in modulus compared to the case with DR2 shown in Fig.2.
Figure 7: The value of $\chi^{2}$ for the back-reaction model with respect to DR1 data as a function of the parameter $A$. The value for $\Lambda$CDM is represented by a star.
Although the best-fit frequency and amplitude of our model show slight variations when using DESI DR2 instead of DR1, and a modest improvement in the $\chi^{2}$ value is observed relative to the $\Lambda$CDM model when using DR2, the overall conclusions remain consistent across both DESI data releases.
Acknowledgements.
We wish to thank Elisa Ferreira for extensive discussions, and E. Colgain for important feedback on the first draft of this paper. M.A.C.A. is supported by Coordenacao de Aperfeicoamento de Pessoal de Nivel Superior (CAPES). L.L.G is supported by research grants from Conselho Nacional de Desenvolvimento Cientıfico e Tecnologico (CNPq), Grant No. 307636/2023-2 and from the Fundacao Carlos Chagas Filho de Amparo a Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/204.598/2024. RB is supported in part by NSERC and the CRC program.